OpenStudy (anonymous):

write the slope intercept equation for the line that passes thru (-2,3) and perpendicular to 2x-9y=-12. show work

OpenStudy (anonymous):

so first you find the slope.. 2x-9y=-12 put this formula into y=....so, -9y=-2x-12 which is y=2/9x+4/3

OpenStudy (anonymous):

slope=2/9

OpenStudy (anonymous):

if it perpendicular you use the reciprocal which is 9/2

OpenStudy (anonymous):

y-y1=m(x-x1) point slope formula ...y-3=9/2(x+2) ...y-3=9/2x +9...= y=9/2x+12

OpenStudy (anonymous):

eliza, the if slope of a line is m, the slope of a line perpendicular to it is negative reciprocal of m, not just the reciprocal of m.

OpenStudy (anonymous):

well just the negative changes

OpenStudy (anonymous):

no, it becomes y = -9/2x -6